OpenStudy (anonymous):

Write the equation of a circle with a center at (0, 8) and a radius of 9.

OpenStudy (whpalmer4):

Do you know the formula for a circle with radius \(r\) and center \((h,k)\)? \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Plug in the values, what do you get?

OpenStudy (anonymous):

what are the values?

OpenStudy (whpalmer4):

the coordinates of the center, and the radius

OpenStudy (anonymous):

I don't know what to plug in to get my answer though

OpenStudy (whpalmer4):

the radius \(r\) = 9 the center \((h,k) = (0,8)\)

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

(x-0)2 + (y-8)2=92

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

If you are going to write exponents, and not bother with the equation editor, at least use the ^ sign to indicate that the following thing is an exponent...otherwise, how am I supposed to know if that last term is "92" or \(9^2\)?

OpenStudy (anonymous):

I thought you would already know so is that the equation?

OpenStudy (whpalmer4):

(x-0)^2 + (y-8)^2 = 9^2 would be acceptable, (x-0)^2+(y-8)^2=81 would probably be better Sure, I know what the equation is, but we're trying to make sure that you know what it is, right? :-)

OpenStudy (anonymous):

alrite

OpenStudy (whpalmer4):

anyhow, that's how you do these problems...any questions about this one?

OpenStudy (whpalmer4):

sometimes you'll see a variation where the problem gives you the two endpoints of a diameter and asks you to find the equation of the circle. Because the diameter is twice the radius, you use the formula for the midpoint of a line to find the midpoint of the diameter, which gives you the center of the circle. Then you use the distance formula to get half the length of the diameter (between the midpoint and either endpoint) so that you have r to fill in that part of the formula.

OpenStudy (anonymous):

ok